Assume that a spherical raindrop evaporates at a rate proportional to its surface area. If its radius
originally is 3 mm, and one-half hour later has been reduced to 2mm, find an expression for the
radius of the raindrop at any time.
1
Expert's answer
2020-02-24T11:34:20-0500

"\frac{dV}{dt}=k*4\pi r^2"

"V=\frac{4}{3}\pi r^3"

"\frac{dV}{dt}=4\pi r^2\frac{dr}{dt}"

So, "\frac{dV}{dt}=4\pi r^2\frac{dr}{dt}=k*4\pi r^2"

or "\frac{dr}{dt}=k"

Integrating, we get: "r=kt+C"

To find k and C we use two conditions: "r(0)=3,\;\; r(0.5)=2"

We have:

"3=k*0+C, \\ \;2=k*0.5+C\;\;or\;\;C=3, k=-2"

And finally, "r=-2t+3" .